# SmallKV: Small Model Assisted Compensation of KV Cache Compression for Efficient LLM Inference

### v1

@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2024b) have emerged with remarkable natural language understanding capabilities and broad application prospects. Despite the advancements, the deployment of LLMs is hindered with significant computational challenges and high GPU memory consumption, particularly when processing long contexts. This issue arises from the intrinsic complexity of their self-attention mechanism, which scales quadratically with the length of the input sequence. Recent developments in reasoning models, exemplified by ChatGPT-o1 (OpenAI, 2024a) and DeepSeek-R1 (DeepSeek-AI, 2025) have exacerbated this issue due to their lengthy reasoning process.
Numerous studies have demonstrated the high degree of sparsity within attention mechanism, leading to the development of various Key-Value cache (KV cache) compression methods such as quantization  (Hooper et al., 2024b; Dong et al., 2024b), eviction (Zhang et al., 2023; Sheng et al., 2023; Yang et al., 2024b), and merging (Kim et al., 2023a; Wang and Xiao, 2024; Nawrot et al., 2024). These methods significantly reduce GPU memory usage and enhance the throughput of inference systems. Our work focuses on eviction-based methods, which identify and retain only the critical tokens to reduce KV cache consumption while minimizing performance degradation in a training-free manner.
@@
A previous study Chen et al. (2021) has revealed a notable similarity in attention patterns between small and large models within the BERT architecture. We further discover the similar observation in the decoder-only architecture models, leading to a novel perspective for addressing the aforementioned limitations. We then propose Small Model Assisted Compensation of KV Cache Compression for Efficient LLM Inference (SmallKV), which introduces a small language model (SLM) to perform saliency shift compensation and marginal information compensation for KV Cache Compression (of LLM). Specifically, the saliency shift compensation mechanism leverages the SLM to maintain global critical information and help identify evicted tokens that may regain significance. Meanwhile, the marginal information compensation identifies tokens with relatively lower attention scores, yet still contribute to model performance. These tokens are less sensitive to the approximation of attention, allowing us to leverage attention scores of SLM for compensation. This results in a hierarchical compression strategy that differentiates and compresses tokens based on their varying levels of importance. It should be noted that SmallKV is compatible with efficient attention implementation such as Flash Attention, which significantly enhances the efficiency of inference. In practical deployment, SmallKV can be used with speculative decoding (Kim et al., 2023b) to speed up LLM inference further.
We conduct a comprehensive evaluation of SmallKV across several benchmarks, including GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2023), MT-bench (Zheng et al., 2023), and Longbench (Bai et al., 2024). The experimental results indicate that SmallKV consistently delivers superior performance, especially under low KV cache budgets. Experiments on different model series (e.g., the Qwen series and LLaMA series) and model sizes (ranging from 7B to 72B) highlight the robustness and generalizability of our approach. Efficiency evaluations show that SmallKV achieves 1.75 - 2.56 times higher throughput compared to previous KV cache compression methods. These experimental results collectively offer compelling evidence of SmallKV’s accuracy and efficiency as a compensation plugin for KV cache compression.
@@ 2 Related Work
The approaches to KV cache compression can be broadly categorized into three classes: eviction, merging, and quantization. Eviction (Oren et al., 2024; Ribar et al., 2023; Lee et al., 2024; Liu et al., 2024d; Zhong et al., 2024; Zhang et al., 2024b) aims to retain only a small set of critical tokens’ KV caches to achieve nearly lossless inference performance. Merging leverages the high angular similarity observed among deep KV caches to reduce both intra-layer (Kim et al., 2023a; Wang and Xiao, 2024; Nawrot et al., 2024; Zhang et al., 2024c; Wan et al., 2024; Wang et al., 2024) and cross-layer (Yang et al., 2024d; Liu et al., 2024b) redundancies. Quantization compresses the data by mapping the original full-precision tensor values to discrete levels and storing them at lower precision, including model weight quantization (Sheng et al., 2023; Yue et al., 2024; Shao et al., 2023) and KV-cache-only quantization (Hooper et al., 2024b; Dong et al., 2024b; Yang et al., 2024c; Dong et al., 2024a; Kang et al., 2024).
@@
Numerous eviction methods have been proposed, which can be categorized into static and dynamic strategies. Static strategies (Ge et al., 2023; Li et al., 2024; Zeng et al., 2024) perform token filtering during prefilling and maintain KV cache of fixed size throughout the subsequent decoding steps (e.g., sliding window attention (Beltagy et al., 2020)). A further approach is to maintain both the initial and the recent tokens (Xiao et al., 2023; Han et al., 2023). Dynamic strategies (Zhang et al., 2023; Yang et al., 2024b; Zhao et al., 2024a; Chen et al., 2024) continuously update the critical KV cache during decoding, while the KV cache of unimportant tokens will be permanently removed or offloaded from the GPU. The core of permanent eviction lies in selecting critical tokens to minimize the damage to model accuracy. H2(Zhang et al., 2023) observed that almost all layers exhibit a sparsity exceeding 95%, indicating that maintaining just 5% of the KV cache based on the accumulative attention scores of each token is sufficient for decoding the same output token at each generation step. PyramidInfer Yang et al. (2024b) employs a pyramid-shaped hierarchical processing approach, where recent tokens are assigned greater weight and the length of KV caches in deeper layers is reduced. Permanent token removal has two main limitations. Irreversible token eviction can degrade model performance in multi-turn dialogue scenarios and long-sequence tasks. Additionally, relying on attention scores prevents the model from adapting to some acceleration techniques (e.g., Flash Attention). Consequently, non-permanent eviction (Xiao et al., 2024; Tang et al., 2024; Zhang et al., 2024a; Hooper et al., 2024a; Liu et al., 2024c) has been proposed by dynamic cache offloading based on indexing. These methods manage and access the KV cache offloaded to multi-tier cache systems at the granularity of chunks or clusters. However, achieving fast and accurate retrieval with high precision remains challenging, while the index construction for retrieval introduces significant decoding latency.
@@ 3.1 Saliency Shift Issue
The dynamic nature of LLM decoding leads to the shift in token saliency, which is the change of the token set with high attention over time. Most existing KV cache eviction strategies, which rely on permanent token removal Zhang et al. (2023); Yang et al. (2024b), are highly susceptible to this issue. To address this, we leverage a small language model (SLM) from the same model series to approximate the LLM’s attention scores. This approach aligns well with speculative decoding (Leviathan et al., 2023; Kim et al., 2023b; Chen et al., 2023; Zhao et al., 2024b; Liu et al., 2024a), where the SLM generates candidate tokens in advance and the LLM verifies them in parallel, thereby allowing the combination of these two appealing approaches for further optimization of inference speed.
@@
Specifically, we quantified the gap caused by saliency shift on the Wikitext-v2 (Merity et al., 2016) dataset using Qwen2-7B in Figure 1 (b). The importance of tokens is measured by accumulative attention scores, following H2OZhang et al. (2023). We measure the discrepancy between sets of important tokens derived from the real drop view and the global view using Jaccard similarity. The relatively low Jaccard similarity values, ranging from 0.55 to 0.77, indicate a significant difference between realistic KV cache eviction methods and the global ground truth, which means a considerable number of globally important tokens are erroneously evicted under the real drop method. In addition, the similarity maintains a clear declining trend as the KV cache budget decreases, indicating that the influence brought by saliency shift is more serious. This is particularly concerning since most current KV cache compression methods claim to operate with a low KV cache budget (e.g., 10%–30%). Therefore, an effective approach to compensate for saliency shift issue is critical and necessary.
@@
We measure the similarity of the attention patterns between Qwen2-0.5B and Qwen2-7B on Wikitext-v2 (Merity et al., 2016) dataset. Figure 1 (c) illustrates one example of the trends of accumulative attention scores between SLM and LLM, as more samples and detailed discussion in Appendix B. For each attention head in LLM, we search the most similar one in the SLM for matching. The quantitative analysis shows that average cosine similarity between Qwen2-0.5B and Qwen2-7B after similarity matching reaches 0.947. The result also visualizes a high degree of consistency, particularly among tokens with high attention scores. This finding indicates a promising way to effectively address the saliency shift issue without introducing significant overhead. We evaluate the effectiveness of saliency shift compensation with 10% KV cache budget across six downstream tasks, as shown in Figure 1 (d). The results demonstrate that using Qwen2-0.5B as the assisted SLM for saliency shift compensation significantly outperforms using only Qwen2-7B without any compensation.
@@ 3.2 Marginal Information Overcompression
Figure 2: Top: The distribution of attention sparsity (blue bar) along with the drop of model accuracy (red line), indicating the necessity of maintaining marginal tokens. Horizontal line (Full) represents the baseline of full KV cache. Bottom: Marginal information compensation by the hierarchical compression that differentiates tokens based on their varying levels of importance.
@@
Observation 2. We measure the accumulative attention scores of each token across all heads of Qwen2.5-14B on the BBH dataset (Suzgun et al., 2023), and statistically rank these scores in Figure 2 to reveal the necessity of maintaining marginal tokens. Consistent with previous research findings, we also observe that only a small fraction of tokens account for the majority of attention scores. The sum of the accumulative attention score for the top 5% of tokens is approximately 7 times greater than that of the 5%-10% range, while the 5%-10% range is only 1.34 times higher than the 10%-15% range, highlighting the dominant position of the top 5% of tokens in attention scores. However, the compression of tokens in range of 5%-15% has caused the accuracy to drop from 0.482 to 0.148 under H2O  method.
@@ 4 SmallKV Method
Unlike previous methods, SmallKV maintains the full cache of SLM $C^{s}_{all}$, and performs eviction for the $i$ cache of LLM based on he $f(i)$-th (full) cache of SLM and the eviction policy $\mathcal{E}(A^{\prime}_{f(i)},C^{s}_{all})$.
@@
where $K$ and $P-K$ denote the numbers of critical tokens and marginal tokens, which are determined by KV cache budget.
@@
The formula shows that for marginal tokens that appear in the $Top(P-K)$ importance, we approximate their attention scores using the corresponding elements from SLM $M_{s}$ to avoid using K cache. This approach allows for retaining important information while evicting the K cache of these tokens, effectively implementing a hierarchical compression policy that differentiates between various levels of token importance. The illustration of attention mechanism in SmallKV is shown in Figure 4.
@@ 5.1 Experiments Setup
Datasets. We comprehensively evaluate the effectiveness of SmallKV in four kinds of scenarios: 1) GSM8K (Cobbe et al., 2021) for mathematical reasoning, 2) BBH (Suzgun et al., 2023) for language understanding, 3) MT-Bench (Zheng et al., 2023) for multi-turn conversation, and 4) Longbench (Bai et al., 2024) for long-context scenario.
@@
H2(Zhang et al., 2023): employs the accumulative attention score as the metric to evict unimportant KV cache.
PyramidInfer (Yang et al., 2024b): identifies that deeper layers exhibit greater redundancy and applies differentiated KV cache budgets between layers.
@@ 5.3 Efficiency Results
Setup. We conduct an efficiency analysis on one NVIDIA A100 GPU. The evaluated model is Qwen2-7B and the SLM used in SmallKV method is Qwen2-0.5B. We use synthetic datasets for testing where all prompts are padded to the same length in the batch and the outputs are also restricted to a fixed length. We consider two common scenarios: 1) Multi-user concurrency: In this scenario, the context is set to a prefill length of 2048 tokens and a decode length of 256 tokens, with a batch size of 64. 2) Extremely long context: the context is set to a prefill length of 16384 tokens and a decode length of 512 tokens, with a batch size of 4. The evaluation metric consists of average time (ms) per output token in decode stage (TPOT), time (s) to first token (TTFT), and end-to-end throughput (token/s). We use Hugging Face Accelerate and H2O as baselines. Hugging Face Accelerate employs two attention implementation methods: eager attention and PyTorch SDPA with Flash Attention kernel (Gschwind et al., ). H2O and SmallKV both employ a KV cache budget of 20%. The times reported in SmallKV include the overhead of the assisted SLM.
@@ 5.4 Ablation Study
Figure 6: Ablation Study of the SmallKV method.
Figure 7: The impact of SLM scaling on SmallKV

### v2

@@ 1 Introduction
Large language models (LLMs) [28] have emerged with remarkable natural language understanding capabilities and broad application prospects. Despite the advancements, the deployment of LLMs is hindered with significant computational challenges and high GPU memory consumption, particularly when processing long contexts. This issue arises from the intrinsic complexity of their self-attention mechanism, which scales quadratically with the length of the input sequence. Recent developments in reasoning models, exemplified by ChatGPT-o1 [29] and DeepSeek-R1 [7] have exacerbated this issue due to their lengthy reasoning process.
Numerous studies have demonstrated the high degree of sparsity within attention mechanism, leading to the development of various Key-Value cache (KV cache) compression methods such as quantization  [15, 9], eviction [51, 33, 43], and merging [17, 38, 27]. These methods significantly reduce GPU memory usage and enhance the throughput of inference systems. Our work focuses on eviction-based methods, which identify and retain only the critical tokens to reduce KV cache consumption while minimizing performance degradation in a training-free manner.
@@
A previous study [4] has revealed a notable similarity in attention patterns between small and large models within the BERT architecture. We further discover the similar observation in the decoder-only architecture models, leading to a novel perspective for addressing the aforementioned limitations. We then propose Small Model Assisted Compensation of KV Cache Compression for Efficient LLM Inference (SmallKV), which introduces a small language model (SLM) to perform saliency shift compensation and marginal information compensation for KV Cache Compression (of LLM). Specifically, the saliency shift compensation mechanism leverages the SLM to maintain global critical information and help identify evicted tokens that may regain significance. Meanwhile, the marginal information compensation identifies tokens with relatively lower attention scores, yet still contribute to model performance. These tokens are less sensitive to the approximation of attention, allowing us to leverage attention scores of SLM for compensation. This results in a hierarchical compression strategy that differentiates and compresses tokens based on their varying levels of importance. It should be noted that SmallKV is compatible with efficient attention implementation such as Flash Attention, which significantly enhances the efficiency of inference. In practical deployment, SmallKV can be used with speculative decoding [18] to speed up LLM inference further.
We conduct a comprehensive evaluation of SmallKV across several benchmarks, including GSM8K [6], BBH [34], MT-bench [54], and Longbench [1]. The experimental results indicate that SmallKV consistently delivers superior performance, especially under low KV cache budgets. Experiments on different model series (e.g., the Qwen series and LLaMA series) and model sizes (ranging from 7B to 72B) highlight the robustness and generalizability of our approach. Efficiency evaluations show that SmallKV achieves 1.75 - 2.56 times higher throughput compared to previous KV cache compression methods. These experimental results collectively offer compelling evidence of SmallKV’s accuracy and efficiency as a compensation plugin for KV cache compression.
@@ 2 Related Work
The approaches to KV cache compression can be broadly categorized into three classes: eviction, merging, and quantization. Eviction [30, 31, 19, 25, 55, 49] aims to retain only a small set of critical tokens’ KV caches to achieve nearly lossless inference performance. Merging leverages the high angular similarity observed among deep KV caches to reduce both intra-layer [17, 38, 27, 50, 37, 39] and cross-layer [45, 23] redundancies. Quantization compresses the data by mapping the original full-precision tensor values to discrete levels and storing them at lower precision, including model weight quantization [33, 46, 32] and KV-cache-only quantization [15, 9, 44, 8, 16].
@@
Numerous eviction methods have been proposed, which can be categorized into static and dynamic strategies. Static strategies [10, 21, 47] perform token filtering during prefilling and maintain KV cache of fixed size throughout the subsequent decoding steps (e.g., sliding window attention [2]). A further approach is to maintain both the initial and the recent tokens [41, 12]. Dynamic strategies [51, 43, 52, 5] continuously update the critical KV cache during decoding, while the KV cache of unimportant tokens will be permanently removed or offloaded from the GPU. The core of permanent eviction lies in selecting critical tokens to minimize the damage to model accuracy. H2[51] observed that almost all layers exhibit a sparsity exceeding 95%, indicating that maintaining just 5% of the KV cache based on the accumulative attention scores of each token is sufficient for decoding the same output token at each generation step. PyramidInfer [43] employs a pyramid-shaped hierarchical processing approach, where recent tokens are assigned greater weight and the length of KV caches in deeper layers is reduced. Permanent token removal has two main limitations. Irreversible token eviction can degrade model performance in multi-turn dialogue scenarios and long-sequence tasks. Additionally, relying on attention scores prevents the model from adapting to some acceleration techniques (e.g., Flash Attention). Consequently, non-permanent eviction [40, 35, 48, 14, 24] has been proposed by dynamic cache offloading based on indexing. These methods manage and access the KV cache offloaded to multi-tier cache systems at the granularity of chunks or clusters. However, achieving fast and accurate retrieval with high precision remains challenging, while the index construction for retrieval introduces significant decoding latency.
Speculative decoding [20, 18, 3, 53] is a widespread technology to accelerate inference. It employs a pair of models, where the smaller one generates candidate tokens in advanced and the larger one verifies them in parallel. By enabling parallel token generation, speculative decoding addresses the inefficiencies inherent in traditional autoregressive approaches, achieving speedups of over three times without compromising the quality of the generated output. This technique has been widely adopted in commercial LLMs [3, 22].
@@ 3.1 Saliency Shift Issue
The dynamic nature of LLM decoding leads to the shift in token saliency, which is the change of the token set with high attention over time. Most existing KV cache eviction strategies, which rely on permanent token removal [51, 43], are highly susceptible to this issue. To address this, we leverage a small language model (SLM) from the same model series to approximate the LLM’s attention scores. This approach aligns well with speculative decoding [20, 18, 3, 53, 22], where the SLM generates candidate tokens in advance and the LLM verifies them in parallel, thereby allowing the combination of these two appealing approaches for further optimization of inference speed.
@@
Specifically, we quantified the gap caused by saliency shift on the Wikitext-v2 [26] dataset using Qwen2-7B in Figure 1 (b). The importance of tokens is measured by accumulative attention scores, following H2O[51]. We measure the discrepancy between sets of important tokens derived from the real drop view and the global view using Jaccard similarity. The relatively low Jaccard similarity values, ranging from 0.55 to 0.77, indicate a significant difference between realistic KV cache eviction methods and the global ground truth, which means a considerable number of globally important tokens are erroneously evicted under the real drop method. In addition, the similarity maintains a clear declining trend as the KV cache budget decreases, indicating that the influence brought by saliency shift is more serious. This is particularly concerning since most current KV cache compression methods claim to operate with a low KV cache budget (e.g., 10%–30%). Therefore, an effective approach to compensate for saliency shift issue is critical and necessary.
@@
We measure the similarity of the attention patterns between Qwen2-0.5B and Qwen2-7B on Wikitext-v2 [26] dataset. Figure 1 (c) illustrates one example of the trends of accumulative attention scores between SLM and LLM, as more samples and detailed discussion in Appendix B. For each attention head in LLM, we search the most similar one in the SLM for matching. The quantitative analysis shows that average cosine similarity between Qwen2-0.5B and Qwen2-7B after similarity matching reaches 0.947. The result also visualizes a high degree of consistency, particularly among tokens with high attention scores. This finding indicates a promising way to effectively address the saliency shift issue without introducing significant overhead. We evaluate the effectiveness of saliency shift compensation with 10% KV cache budget across six downstream tasks, as shown in Figure 1 (d). The results demonstrate that using Qwen2-0.5B as the assisted SLM for saliency shift compensation significantly outperforms using only Qwen2-7B without any compensation.
@@ 3.2 Marginal Information Overcompression
Figure 2: Left: The distribution of attention sparsity (blue bar) along with the drop of model accuracy (red line), indicating the necessity of maintaining marginal tokens. Horizontal line (Full) represents the baseline of full KV cache. Right: Marginal information compensation by the hierarchical compression that differentiates tokens based on their varying levels of importance.
Observation 2. We measure the accumulative attention scores of each token across all heads of Qwen2.5-14B on the BBH dataset [34], and statistically rank these scores in Figure 2 to reveal the necessity of maintaining marginal tokens. Consistent with previous research findings, we also observe that only a small fraction of tokens account for the majority of attention scores. The sum of the accumulative attention score for the top 5% of tokens is approximately 7 times greater than that of the 5%-10% range, while the 5%-10% range is only 1.34 times higher than the 10%-15% range, highlighting the dominant position of the top 5% of tokens in attention scores. However, the compression of tokens in range of 5%-15% has caused the accuracy to drop from 0.482 to 0.148 under H2O  method.
@@ 4 SmallKV Method
Unlike previous methods, SmallKV maintains the full cache of SLM $C^{s}_{all}$, and performs eviction for the $i$ cache of LLM based on the $f(i)$-th (full) cache of SLM and the eviction policy $\mathcal{E}(A^{\prime}_{f(i)},C^{s}_{all})$.
@@
where $K$ and PKP-K denote the numbers of critical tokens and marginal tokens, which are determined by KV cache budget.
@@
The formula shows that for marginal tokens that appear in the Top(PK)Top(P-K) importance, we approximate their attention scores using the corresponding elements from SLM $M_{s}$ to avoid using K cache. This approach allows for retaining important information while evicting the K cache of these tokens, effectively implementing a hierarchical compression policy that differentiates between various levels of token importance. The illustration of attention mechanism in SmallKV is shown in Figure 4.
@@ 5.1 Experiments Setup
Datasets. We comprehensively evaluate the effectiveness of SmallKV in four kinds of scenarios: 1) GSM8K [6] for mathematical reasoning, 2) BBH [34] for language understanding, 3) MT-Bench [54] for multi-turn conversation, and 4) Longbench [1] for long-context scenario.
@@
H2[51]: employs the accumulative attention score as the metric to evict unimportant KV cache.
PyramidInfer [43]: identifies that deeper layers exhibit greater redundancy and applies differentiated KV cache budgets between layers.
@@ 5.3 Efficiency Results
Setup. We conduct an efficiency analysis on one NVIDIA A100 GPU. The evaluated model is Qwen2-7B and the SLM used in SmallKV method is Qwen2-0.5B. We use synthetic datasets for testing where all prompts are padded to the same length in the batch and the outputs are also restricted to a fixed length. We consider two common scenarios: 1) Multi-user concurrency: In this scenario, the context is set to a prefill length of 2048 tokens and a decode length of 256 tokens, with a batch size of 64. 2) Extremely long context: the context is set to a prefill length of 16384 tokens and a decode length of 512 tokens, with a batch size of 4. The evaluation metric consists of average time (ms) per output token in decode stage (TPOT), time (s) to first token (TTFT), and end-to-end throughput (token/s). We use Hugging Face Accelerate and H2O as baselines. Hugging Face Accelerate employs two attention implementation methods: eager attention and PyTorch SDPA with Flash Attention kernel [11]. H2O and SmallKV both employ a KV cache budget of 20%. The times reported in SmallKV include the overhead of the assisted SLM.
@@ 5.4 Ablation Study
Figure 6: Ablation Study of SmallKV method.
Figure 7: The impact of SLM scaling on SmallKV.
@@ A.2 Computational overhead
A.3 Quantitative Analysis and Optimization of SLM
We quantify the theoretical maximum latency and memory overhead of the model combinations used in main experiments, as shown in the Table  4. The values ’Ratio’ represent the additional overhead introduced by SLM relative to LLM (e.g., for combination Qwen2-0.5B & 7B, the KV cache of SLM is 1/4.67 that of the LLM).
Table 4: Quantitative analysis of SLM relative to LLM in latency and KV cache usage
Ratio between SLM and LLM Qwen2-0.5B & Qwen2-7B Qwen2.5-0.5B & Qwen2.5-14B Qwen2-7B & Qwen2-72B LLAMA 3.2-1B & LLAMA 3.1-8B
Latency 1/3.57 1/6.52 1/4.47 1/3.19
KV cache 1/4.67 1/16.0 1/5.71 1/4.0
We propose three deployment-level optimizations to effectively reduce the computational and memory overhead of the SLM, enhancing the practicality of SmallKV: (1) integration with speculative decoding, (2) compression of the KV cache for the SLM, and (3) layer skipping (i.e., early stopping) during the attention mapping process.
(1) The integration of speculative decoding can eliminate the forward pass of the SLM, significantly reduce latency introduced by the SLM. Current speculative decoding algorithms yield 3-5x speedups, which would further enhance performance.
(2) The high attention sparsity inherent in SLM enables near-lossless reduction in memory usage through SLM KV cache compression. As shown in the Table  5, we fix the KV cache budget of the LLM (Qwen2-7B) at 20% and increase the compression ratio of the SLM (Qwen2-0.5B) to evaluate the LLM’s accuracy on BBH (SmallKV ACC = 0.436 at 20% KV cache budget). Under the same memory constraint (for SmallKV we count both SLM and LLM cache), the performance of SmallKV (ACC = 0.462 at 40% KV cache budget for SLM and 20% for LLM) outperforms H2O (ACC = 0.38 at 30% KV cache budget). Note that while SLM compression may slightly impair saliency shift compensation, the trade-off remains favorable in terms of overall efficiency.
(3) In an independently deployed scenario, the SLM is not required to generate draft tokens, enabling the adoption of an early layer stopping strategy. This approach significantly reduces both latency and memory overhead. We fix the LLM’s KV cache budget at 20% and progressively reduce the number of layers utilized in the SLM for attention mapping. The results shown in Table  6 indicate that stopping at layer 20 (83% of total layers) incurs no loss in accuracy (0.436). Furthermore, even when halting at layer 16 (67%), the accuracy of SmallKV remains superior to that of H2O.
Table 5: Accuracy of LLM under different KV cache budgets of SLM
KV Cache Budget of SLM 20% 40% 60% 80% Full
Accuracy (BBH) 0.2873 0.4416 0.4624 0.4642 0.4359
Table 6: Accuracy of LLM under different stop layers of SLM attention mapping
Stop Layer of SLM 24 (Full) 20 18 16 12 H3O
Accuracy (BBH) 0.436 0.436 0.390 0.359 0.335 0.318
Based on the implementation optimization of (2) and (3), the SLM’s actual KV cache usage is calculated as: (1 / 4.67) $\times$ 40% (SLM compression) $\times$ (20 / 24) (early stopping) $\approx$ 7.14% of the full LLM KV cache.